# SPO: Sequential Monte Carlo Policy Optimisation

### v2

@@ G.4 Connection to Mirror Descent Guided Policy Search
3:    S-step: $\pi_{\theta}\leftarrow\arg\min_{\theta}\sum_{i}D_{\text{KL}}(p_{i}(\tau)\parallel\pi_{\theta}(\tau))$ (via supervised learning)
@@
Comparing MD-GPS to EM, the C-step is equivalent to the E-step and S-step equivalent to the M-step, with the loss $\ell$ being the expected discounted sum of returns over trajectories in EM. MD-GPS assumes that the S-step can perfectly minimise the objective in the case of linear dynamics and quadritic costs resulting in exact mirror descent. In practice, most applications of MD-GPS will not satisfy such constraints and so are akin to approximate mirror descent. However as long as a constraint between the local and global policy can be enforced in terms of KL various bounds on cost of the global policy can be constructed, see Montgomery and Levine [54] for further details.

### v3

@@ G.6 Connection to Mirror Descent Guided Policy Search
3:    S-step: $\pi_{\theta}\leftarrow\arg\min_{\theta}\sum_{i}D_{\text{KL}}(p_{i}(\tau)\parallel\pi_{\theta}(\tau))$
@@
Comparing MD-GPS to EM, the C-step is equivalent to the E-step and S-step equivalent to the M-step, with the loss $\ell$ being the negative of the expected discounted sum of returns over trajectories. MD-GPS assumes that the S-step can perfectly minimise the objective in the case of linear dynamics and quadratic costs [48], resulting in exact mirror descent. In practice, most applications of MD-GPS will not satisfy such constraints and so are akin to approximate mirror descent. However, as long as a constraint between the local and global policy can be enforced in terms of KL, various bounds on cost of the global policy can be constructed, see Montgomery and Levine [58] for further details.